### v1

@@ Abstract
scientific machine learning, nonlinear preconditioning, Schwarz methods, domain-decomposition
{MSCcodes}
90C30, 90C26, 90C06, 65M55, 68T07
@@ 1 Introduction
Both, the physics informed and operator learning, approaches are commonly trained using gradient-based optimizers, e.g., Adam [36] or LBFGS [51]. However, the computational cost of these optimizers can be exorbitant due to the non-convexity and ill-conditioning of the underlying optimization problem. This is especially the case when a highly accurate solution is required. In order to reduce the training time, the parallelization of the DNN training process has gained significant attention in the literature; see, for example [5, 13]. Two of the most popular parallel paradigms are data-parallel and model-parallel approaches. Data-parallel approaches [45] split the training data among multiple processing devices. Each device then evaluates the loss/gradient for a given portion of data, and the average value is used to update the network parameters. In contrast, model-parallel approaches aim at splitting the parameters of the DNN model among multiple devices, see for example pipelining techniques [30, 58] or tensor-parallel approaches [5, 61].
@@
  • We introduce overlapping layer-wise decomposition of DNNs, effectively improving the convergence of non-overlapping layer-wise preconditioners.
  • A novel subdomain-wise synchronization strategy is proposed to recombine corrections obtained from parallel subdomain training. This expands upon commonly used methodologies in numerical analysis, which combine subdomain corrections using the weighted averaging technique, see for example literature on the (nonlinear) restricted ASM [9, 10].
  • We incorporate a coarse-level training step, which enables the global communication between subdomains while maintaining the feed-forward nature of feature flow in the DNNs.
  • The applicability of the single-level and eventually two-level layer-wise preconditioners is extended beyond PINNs to DeepONets.
  • Through numerical experiments, we demonstrate that the proposed two-level preconditioner yields more accurate DNN models, while also improving the convergence rate of the standard LBFGS optimizer and the baseline single-level preconditioner [40].
This paper is organized as follows. Section 2 provides an introduction to scientific machine learning, with a particular focus on PINNs and DeepONets. In Section 3, we propose a novel two-level overlapping additive Schwarz preconditioner for training DNNs. Section 4 presents a set of benchmark problems and discusses the implementation details. Finally, in Section 5, we demonstrate the numerical performance of the proposed preconditioner. The conclusion and future work are discussed in Section 6.
@@ 2 Scientific machine learning
(1) $\begin{split}\pazocal{A}(u(\boldsymbol{x},\boldsymbol{\eta}),\boldsymbol{\eta})&=f(\boldsymbol{x},\boldsymbol{\eta}),\quad\ \boldsymbol{x}\in\Omega,\\ \pazocal{B}^{k}(u(\boldsymbol{x},\boldsymbol{\eta}),\boldsymbol{\eta})&=g^{k}(\boldsymbol{x},\boldsymbol{\eta}),\quad\boldsymbol{x}\in\Gamma^{k}\subseteq\partial\Omega,\quad\text{ for }k=1,2,\cdots,n_{\Gamma},\end{split}$
@@ 2.1 Physics informed neural networks
(2) $\boldsymbol{\theta}^{\ast}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\mathbb{R}^{n}}\pazocal{L}(\boldsymbol{\theta}):=w_{\mathrm{int}}\pazocal{L}_{\mathrm{int}}(\boldsymbol{\theta})+w_{\mathrm{bc}}\pazocal{L}_{\mathrm{bc}}(\boldsymbol{\theta}),$
@@
(3) $\begin{split}\pazocal{L}_{\mathrm{int}}(\boldsymbol{\theta})&:=\|\pazocal{A}(u_{\boldsymbol{\theta}})-f\|^{2}_{L^{2}(\Omega)}=\int_{\Omega}(\pazocal{A}(u_{\boldsymbol{\theta}}(\boldsymbol{x}))-f(\boldsymbol{x}))^{2}d\boldsymbol{x},\\ \pazocal{L}_{\mathrm{bc}}(\boldsymbol{\theta})&:=\|\pazocal{B}(u_{\boldsymbol{\theta}})-g\|^{2}_{L^{2}(\partial\Omega)}=\int_{\partial\Omega}(\pazocal{B}(u_{\boldsymbol{\theta}}(\boldsymbol{x}))-g(\boldsymbol{x}))^{2}d\boldsymbol{x}.\end{split}$
@@
$\begin{split}\pazocal{L}_{\mathrm{int}}(\boldsymbol{\theta})&=\frac{1}{|\pazocal{D}_{\mathrm{int}}|}\sum_{\boldsymbol{x}_{j}\in\pazocal{D}_{\mathrm{int}}}(\pazocal{A}(u_{\boldsymbol{\theta}}(\boldsymbol{x}_{j}))-f(\boldsymbol{x}_{j}))^{2},\\ \pazocal{L}_{\mathrm{bc}}(\boldsymbol{\theta})&=\sum_{k=1}^{n_{\Gamma}}\left(\frac{1}{|\pazocal{D}^{k}_{\mathrm{bc}}|}\sum_{\boldsymbol{x}^{k}_{j}\in\pazocal{D}^{k}_{\mathrm{bc}}}(\pazocal{B}(u_{\boldsymbol{\theta}}(\boldsymbol{x}^{k}_{j}))-g(\boldsymbol{x}^{k}_{j}))^{2}\right).\end{split}$
The weights $w_{\mathrm{int}}$ and $w_{\mathrm{bc}}$ in (2) are used to balance the interior and boundary loss terms. In practice, the quality of the DNN approximation and the success of the training process depends on the appropriate selection of these weights [64]. Consequently, many approaches for manual or automatic selection of weights have been proposed in the literature; see, for instance, NTK-based weighting strategies [68] or residual-based attention mechanism [1]. In this work, we opt to avoid selecting $w_{\mathrm{int}},w_{\mathrm{bc}}$ by directly incorporating the boundary conditions into the network architecture. Following [54, 64], we modify the network output as
(4) $\tilde{u}_{\boldsymbol{\theta}}(\boldsymbol{x}):=\sum_{k=1}^{n_{\Gamma}}g^{k}(\boldsymbol{x})+\ell(\boldsymbol{x})u_{\boldsymbol{\theta}}(\boldsymbol{x}),\quad\text{ where }\ell(\boldsymbol{x}):=\prod_{k=1}^{n_{\Gamma}}(1-(1-\ell^{k}(\boldsymbol{x}))).$
@@
$\ell^{k}(\boldsymbol{x})=\begin{cases}0,&\boldsymbol{x}\in\Gamma^{k},\\ 1,&\boldsymbol{x}\in\Gamma^{j},\quad j\in\{1,2,\cdots,n_{\Gamma}\},\quad\text{for }k\neq j,\\ t\in(0,1),&\text{otherwise}.\end{cases}$
Using $\tilde{u}_{\boldsymbol{\theta}}$ as defined in (4), we can now reformulate the minimization problem (2) as
(5) $\boldsymbol{\theta}^{\ast}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\mathbb{R}^{n}}\pazocal{L}(\boldsymbol{\theta}):=\frac{1}{|\pazocal{D}_{\mathrm{int}}|}\sum_{\boldsymbol{x}_{j}\in\pazocal{D}_{\mathrm{int}}}(\pazocal{A}(\tilde{u}_{\boldsymbol{\theta}}(\boldsymbol{x}_{j}))-f(\boldsymbol{x}_{j}))^{2}.$
@@ 2.2 Operator learning
The architecture of DON consists of branch and trunk networks. The branch network $B\colon\pazocal{Y}_{m}\to\mathbb{R}^{p}$ provides a set of $p$ coefficients for a given discretized function ${\boldsymbol{y}\in\pazocal{Y}_{m}}$. Here, $\pazocal{Y}_{m}$ is a $m$-dimensional space that is assumed to satisfy the following condition:
$\forall y\in\pazocal{Y},\ \exists\boldsymbol{y}_{m}\in\pazocal{Y}_{m}\text{ such that }\boldsymbol{y}_{m}\to y\text{ as }m\to\infty.$
In addition, the trunk network $T\colon\Omega\to\mathbb{R}^{p}$ is used to encode the computational domain $\Omega$. Therefore, it takes as an input a coordinate $\boldsymbol{\xi}\in\Omega$ and returns a set of $p$ basis functions.
@@
(6) $\pazocal{G}_{\boldsymbol{\theta}}(\boldsymbol{y})(\boldsymbol{\xi}):=\sum_{k=1}^{p}B_{k}(\boldsymbol{y})T_{k}(\boldsymbol{\xi}),\quad\boldsymbol{y}\in\pazocal{Y}_{m},\boldsymbol{\xi}\in\Omega,$
to approximate a solution of the parametric PDE at a point $\boldsymbol{\xi}$ for a given function $\boldsymbol{y}$. Figure 1 illustrates a prototypical DON architecture. Note that we are free to choose the architecture of branch and trunk networks. For example, the choice of the branch network can guided by the structure of the input $\boldsymbol{y}$, e.g., it can be a fully connected network in case of a scalar inputs, or a convolutional network in case of multidimensional discretized functions. Since the vector of coordinates $\boldsymbol{\xi}\in\mathbb{R}^{d}$ is usually low dimensional, fully connected networks are commonly used in place of the trunk network. Alternatively, one can utilize the precomputed basis functions, e.g., obtained by means of the proper orthogonal decomposition (POD) performed on the training data [53] or by means of parameterizing the integral kernel in the Fourier space [2].
@@
consisting of $n_{s}$ samples. Each triple $(\boldsymbol{y}_{j},\bar{\boldsymbol{\xi}}_{j},\boldsymbol{u}_{j})$ consists of the discretized function ${\boldsymbol{y}_{j}\in\mathbb{R}^{m}}$, sampled from $\pazocal{Y}_{m}$. Moreover, $\bar{\boldsymbol{\xi}}_{j}=[\boldsymbol{\xi}_{j,1},\cdots,\boldsymbol{\xi}_{j,n_{c}}]^{T}\in\mathbb{R}^{n_{c}\times d}$ represents a set of $n_{c}$ collocation points and $\boldsymbol{u}_{j}\in\mathbb{R}^{n_{c}}$ is a target solution evaluated at collocation points constituting $\bar{\boldsymbol{\xi}}_{j}$.
@@
(7) $\boldsymbol{\theta}^{\ast}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\mathbb{R}^{n}}\pazocal{L}(\boldsymbol{\theta}):=\frac{1}{n_{s}n_{c}}\sum_{j=1}^{n_{s}}\sum_{t=1}^{n_{c}}\left(\pazocal{G}_{\boldsymbol{\theta}}(\boldsymbol{y}_{j})(\boldsymbol{\xi}_{j,t})-\boldsymbol{u}_{j}\right)^{2}.$
@@
Figure 1: Left: A sketch of DON architecture. The trunk can be represented by the DNN or by the POD basis obtained from the training data. The symbol $\boldsymbol{\times}$ denotes the inner product between the output of the branch and trunk networks. Right: A relative error obtained by using the standard LBFGS optimizer and the proposed TL-LBFGS as a function of increasing number of layers. The experiment is performed for the AC example, specified in Section 4.1.
@@ 3 Two-level overlapping additive Schwarz preconditioner
(8) $\min_{\boldsymbol{\theta}}\pazocal{L}(\boldsymbol{\theta}),$
where the loss $\pazocal{L}$ is as defined in (5) or (7) for PINN or DON, respectively. The minimization problem (8) is traditionally solved using gradient-based optimizers such as Adam [36] or LBFGS [51]. However, the convergence of these optimizers tends to deteriorate as problem stiffness and ill-conditioning increase [66], resulting in slow training and unsatisfactory accuracy of the resulting PINN/DON models. For instance, Fig. 1 illustrates how increasing the number of layers in a PINN model exacerbates the ill-conditioning and non-convexity of the underlying minimization problem. This, in turn, presents significant challenges for the standard LBFGS method to converge, leading to poorer accuracy of the PINN as the representation capacity of the DNN increases.
To overcome these difficulties, the single-level Schwarz preconditioning strategy, which utilizes the layer-wise decomposition of the DNN, was introduced in [40]. In this work, we propose enhancing the performance of this strategy by introducing overlap, a coarse-level correction step, and a subdomain-based recombination strategy. The culmination of these techniques results in a novel two-level overlapping Schwarz preconditioner, the performance of which does not deteriorate with increasing network depth, as shown in Fig. 1.
@@
(9) $\nabla\pazocal{L}(\boldsymbol{\theta})=0.$
@@
(10) $\pazocal{F}(\boldsymbol{\theta}):=\nabla\pazocal{L}(\pazocal{M}(\boldsymbol{\theta}))=0,$
where the preconditioner $\pazocal{M}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is designed such that (10) has more balanced nonlinearities than (9).
In this work, we solve the nonlinear system (10) iteratively. Thus, on each $k$-th iteration, we start from a given iterate $\boldsymbol{\theta}^{(k)}$, and apply the preconditioner $\pazocal{M}$. As a result, we obtain the smoothed iterate $\boldsymbol{\theta}^{(k+1/2)}:=\pazocal{M}(\boldsymbol{\theta}^{(k)})$, which is required to reduce the loss, i.e., it has to satisfy the following condition:

### v2

@@ Abstract
Scientific machine learning , Nonlinear preconditioning , Schwarz methods , Domain decomposition
journal: Computer Methods in Applied Mechanics and Engineering
[…]
@@ 1 Introduction
Both, physics informed and operator learning, approaches are commonly trained using gradient-based optimizers, e.g., Adam [11] or LBFGS [12]. However, the computational cost of these optimizers can be exorbitant due to the non-convexity and ill-conditioning of the underlying optimization problem. To address this problem, there has been research on improving these optimizers [13, 14, 15]. This is especially the case when a highly accurate solution is required. In order to reduce the training time, the parallelization of the DNN training process has gained significant attention in the literature; see, for example [16, 17]. Two of the most popular parallel paradigms are data-parallel and model-parallel approaches. Data-parallel approaches [18] split the training data among multiple processing devices. Each device then evaluates the loss/gradient for a given portion of data, and the average value is used to update the network parameters. In contrast, model-parallel approaches aim at splitting the parameters of the DNN model among multiple devices, see for example pipelining techniques [19, 20] or tensor-parallel approaches [16, 21].
@@
  • We introduce overlapping layer-wise decomposition of DNNs, effectively improving the convergence of non-overlapping layer-wise preconditioners. Motivated by domain-decomposition (DD) approaches in scientific computing, we enforce overlap between the parameters of neighboring layers/subdomains.
  • A novel subdomain-wise synchronization strategy is proposed to recombine corrections obtained from parallel subdomain training. This expands upon commonly used methodologies in numerical analysis, which combine subdomain corrections using the weighted averaging technique, see for example literature on the (nonlinear) restricted additive Schwarz method [37, 42]. In the context of the layer-wise DD methods considered in this work, introducing a subdomain-wise synchronization strategy improves the convergence speed of the overlapping DD method as the size of the overlap increases.
  • We incorporate a coarse-level training step that enables the global communication between subdomains while maintaining the feed-forward nature of feature flow in the DNNs. This leads to significantly faster convergence compared to using single-level domain decomposition methods.
  • The applicability of the single-level and eventually two-level layer-wise preconditioners is extended beyond PINNs to DeepONets, which demonstrate the applicability of the proposed method to wider range of machine-learning problems.
  • Through a series of numerical experiments, we demonstrate that the proposed two-level preconditioner yields more accurate DNN models, while also improving the convergence rate of the standard LBFGS optimizer and the baseline single-level preconditioner [30].
This paper is organized as follows. Section˜2 provides an introduction to scientific machine learning, with a particular focus on PINNs and DeepONets. In Section˜3, we propose a novel two-level overlapping additive Schwarz preconditioner for training DNNs. Section˜4 presents a set of benchmark problems and discusses the implementation details. Finally, in Section˜5, we demonstrate the numerical performance of the proposed preconditioner. The conclusion and future work are discussed in Section˜7.
@@ 2 Scientific machine learning
$\begin{split}\pazocal{A}(u(\boldsymbol{x},\boldsymbol{\eta}),\boldsymbol{\eta})&=f(\boldsymbol{x},\boldsymbol{\eta}),\quad\ \boldsymbol{x}\in\Omega,\\ \pazocal{B}^{k}(u(\boldsymbol{x},\boldsymbol{\eta}),\boldsymbol{\eta})&=g^{k}(\boldsymbol{x},\boldsymbol{\eta}),\quad\boldsymbol{x}\in\Gamma^{k}\subseteq\partial\Omega,\quad\text{ for }k=1,2,\ldots,n_{\Gamma},\end{split}$ (1)
@@ 2.1 Physics informed neural networks
$\boldsymbol{\theta}^{\ast}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\mathbb{R}^{n}}\pazocal{L}(\boldsymbol{\theta}):=w_{\mathrm{int}}\pazocal{L}_{\mathrm{int}}(\boldsymbol{\theta})+w_{\mathrm{bc}}\pazocal{L}_{\mathrm{bc}}(\boldsymbol{\theta}),$ (2)
@@
$\begin{split}\pazocal{L}_{\mathrm{int}}(\boldsymbol{\theta})&:=\|\pazocal{A}(u_{\boldsymbol{\theta}})-f\|^{2}_{L^{2}(\Omega)}=\int_{\Omega}(\pazocal{A}(u_{\boldsymbol{\theta}}(\boldsymbol{x}))-f(\boldsymbol{x}))^{2}d\boldsymbol{x},\\ \pazocal{L}_{\mathrm{bc}}(\boldsymbol{\theta})&:=\|\pazocal{B}(u_{\boldsymbol{\theta}})-g\|^{2}_{L^{2}(\partial\Omega)}=\int_{\partial\Omega}(\pazocal{B}(u_{\boldsymbol{\theta}}(\boldsymbol{x}))-g(\boldsymbol{x}))^{2}d\boldsymbol{x}.\end{split}$ (3)
@@
$\begin{split}\pazocal{L}_{\mathrm{int}}(\boldsymbol{\theta})&=\frac{1}{|\pazocal{D}_{\mathrm{int}}|}\sum_{\boldsymbol{x}_{j}\in\pazocal{D}_{\mathrm{int}}}(\pazocal{A}(u_{\boldsymbol{\theta}}(\boldsymbol{x}_{j}))-f(\boldsymbol{x}_{j}))^{2},\\ \pazocal{L}_{\mathrm{bc}}(\boldsymbol{\theta})&=\sum_{k=1}^{n_{\Gamma}}\left(\frac{1}{|\pazocal{D}^{k}_{\mathrm{bc}}|}\sum_{\boldsymbol{x}^{k}_{j}\in\pazocal{D}^{k}_{\mathrm{bc}}}(\pazocal{B}^{k}(u_{\boldsymbol{\theta}}(\boldsymbol{x}^{k}_{j}))-g^{k}(\boldsymbol{x}^{k}_{j}))^{2}\right).\end{split}$
The weights $w_{\mathrm{int}}$ and $w_{\mathrm{bc}}$ in (2) are used to balance the interior and boundary loss terms. In practice, the quality of the DNN approximation and the success of the training process depends on the appropriate selection of these weights [44]. Consequently, many approaches for manual or automatic selection of weights have been proposed in the literature; see, for instance, NTK-based weighting strategies [45] or residual-based attention mechanism [46]. In this work, Dirichlet boundary conditions are directly embedded into the network architecture, thereby avoiding explicit selection of $w_{\mathrm{int}}$ and $w_{\mathrm{bc}}$. Following [44, 47], we modify the network output as
$\tilde{u}_{\boldsymbol{\theta}}(\boldsymbol{x}):=\sum_{k=1}^{n_{\Gamma}}g^{k}(\boldsymbol{x})+\ell(\boldsymbol{x})u_{\boldsymbol{\theta}}(\boldsymbol{x}),\quad\text{ where }\ell(\boldsymbol{x}):=\prod_{k=1}^{n_{\Gamma}}\ell^{k}(\boldsymbol{x}).$ (4)
@@
$\ell^{k}(\boldsymbol{x})=\begin{cases}0,&\boldsymbol{x}\in\Gamma^{k},\\ 1,&\boldsymbol{x}\in\Gamma^{j},\quad j\in\{1,2,\ldots,n_{\Gamma}\},\quad\text{ for }k\neq j,\\ t\in(0,1),&\text{otherwise}.\end{cases}$
Note that $t$ is determined so that $\ell^{k}(\boldsymbol{x})$ is a smooth function. For example, when $\Omega=(0,1)^{2}$ and $\partial\Omega=\cup_{k=1}^{n_{\Gamma}}\Gamma^{k}$, $\ell(\boldsymbol{x})=\ell(x_{1},x_{2})=x_{1}(1-x_{1})x_{2}(1-x_{2})$. Using $\tilde{u}_{\boldsymbol{\theta}}$ as defined in (4), we can now reformulate the minimization problem (2) as
$\boldsymbol{\theta}^{\ast}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\mathbb{R}^{n}}\pazocal{L}(\boldsymbol{\theta}):=\frac{1}{|\pazocal{D}_{\mathrm{int}}|}\sum_{\boldsymbol{x}_{j}\in\pazocal{D}_{\mathrm{int}}}(\pazocal{A}(\tilde{u}_{\boldsymbol{\theta}}(\boldsymbol{x}_{j}))-f(\boldsymbol{x}_{j}))^{2}.$ (5)
@@ 2.2 Operator learning
The architecture of DON consists of branch and trunk networks. The branch network $B\colon\pazocal{Y}_{m}\to\mathbb{R}^{p}$ provides a set of $p$ coefficients for a given discretized function $\boldsymbol{y}^{m}\in\pazocal{Y}_{m}$}. Here, $\pazocal{Y}_{m}$ is a $m$-dimensional space that is assumed to satisfy the following condition:
$\forall y\in\pazocal{Y},\ \exists\boldsymbol{y}^{m}\in\pazocal{Y}_{m}\subseteq\pazocal{Y}\text{ such that }\boldsymbol{y}^{m}\to y\text{ as }m\to\infty.$
For simplicity, we write $\boldsymbol{y}$ in place of $\boldsymbol{y}^{m}$, as $m$ is a fixed and determines the dimension of the discrete space. In addition, the trunk network $T\colon\Omega\to\mathbb{R}^{p}$ is used to encode the computational domain $\Omega$. Therefore, it takes as an input a coordinate $\boldsymbol{\xi}\in\Omega$ and returns a set of $p$ basis functions.
@@
$\pazocal{G}_{\boldsymbol{\theta}}(\boldsymbol{y})(\boldsymbol{\xi}):=\sum_{k=1}^{p}B_{k}(\boldsymbol{y})T_{k}(\boldsymbol{\xi}),\quad\boldsymbol{y}\in\pazocal{Y}_{m},\boldsymbol{\xi}\in\Omega,$ (6)
to approximate a solution of the parametric PDE at a point $\boldsymbol{\xi}$ for a given function $\boldsymbol{y}$. Figure 1 illustrates a prototypical DON architecture. Note that we are free to choose the architecture of branch and trunk networks. For example, the choice of the branch network can be guided by the structure of the input $\boldsymbol{y}$, e.g., it can be a fully connected network in case of a scalar input, or a convolutional network in case of multidimensional discretized functions. Since the vector of coordinates $\boldsymbol{\xi}\in\mathbb{R}^{d}$ is usually low dimensional, fully connected networks are commonly used in place of the trunk network. Alternatively, one can utilize the precomputed basis functions, e.g., obtained by means of the proper orthogonal decomposition (POD) performed on the training data [48] or by means of parameterizing the integral kernel in the Fourier space [49].
@@
consisting of $n_{s}$ samples. Each triple $(\boldsymbol{y}_{j},\bar{\boldsymbol{\xi}}_{j},\boldsymbol{u}_{j})$ consists of the discretized function $\boldsymbol{y}_{j}\in\mathbb{R}^{m}$}, sampled from $\pazocal{Y}_{m}$. Moreover, $\bar{\boldsymbol{\xi}}_{j}=[\boldsymbol{\xi}_{j,1},\ldots,\boldsymbol{\xi}_{j,n_{c}}]^{T}\in\mathbb{R}^{n_{c}\times d}$ represents a set of $n_{c}$ collocation points and $\boldsymbol{u}_{j}\in\mathbb{R}^{n_{c}}$ is a target solution evaluated at collocation points constituting $\bar{\boldsymbol{\xi}}_{j}$. Note that the target solution is typically obtained using high-fidelity discretization techniques, such as finite element or finite difference methods.
@@
$\boldsymbol{\theta}^{\ast}=\operatorname*{arg\,min}_{\boldsymbol{\theta}\in\mathbb{R}^{n}}\pazocal{L}(\boldsymbol{\theta}):=\frac{1}{n_{s}n_{c}}\sum_{j=1}^{n_{s}}\sum_{t=1}^{n_{c}}\left(\pazocal{G}_{\boldsymbol{\theta}}(\boldsymbol{y}_{j})(\boldsymbol{\xi}_{j,t})-\boldsymbol{u}_{j}\right)^{2}.$ (7)
@@
[…]
@@ 3 Two-level overlapping additive Schwarz preconditioner
$\min_{\boldsymbol{\theta}}\pazocal{L}(\boldsymbol{\theta}),$ (8)
where the loss $\pazocal{L}$ is as defined in (5) or (7) for PINN or DON, respectively. The minimization problem (8) is traditionally solved using gradient-based optimizers such as Adam [11] or LBFGS [12]. However, the convergence of these optimizers tends to deteriorate as problem stiffness and ill-conditioning increase [50], resulting in slow training and unsatisfactory accuracy of the resulting PINN/DON models. For instance, Figure˜1 illustrates that increasing the number of layers in a PINN model exacerbates the relative error when using a standard optimizer. As the depth of the network grows, the ill-conditioning and non-convexity of the underlying minimization problem become more significant. This, in turn, presents significant challenges for the standard LBFGS method to converge, leading to poorer accuracy of the PINN as the representation capacity of the DNN increases.
In scientific computing, Schwarz methods [24, 51] provide an efficient framework for solving large-scale problems by decomposing the solution space into smaller subspaces, within which the problem can be solved more effectively. Emulating the classical additive Schwarz methods, the single-level Schwarz preconditioning strategy, which utilizes the layer-wise decomposition of the DNN, was introduced in [30]. In this work, we propose enhancing the performance of this strategy by introducing overlap, a coarse-level correction step, and a subdomain-based recombination strategy. The culmination of these techniques results in a novel two-level overlapping Schwarz preconditioner, the performance of which does not deteriorate with increasing network depth, as shown in Figure˜1.
@@
$\nabla\pazocal{L}(\boldsymbol{\theta})=0.$ (9)
@@
$\pazocal{F}(\boldsymbol{\theta}):=\nabla\pazocal{L}(\pazocal{M}(\boldsymbol{\theta}))=0,$ (10)
where the preconditioner $\pazocal{M}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is designed such that (10) has more balanced nonlinearities than (9). In practice, this means that the nonlinear terms are of comparable magnitude both to each other and to the linear terms, thereby avoiding dominance by any single contribution.
In this work, we solve the nonlinear system (10) iteratively. Thus, on each $k$-th iteration, we start from a given iterate $\boldsymbol{\theta}^{(k)}$, and apply the preconditioner $\pazocal{M}$. As a result, we obtain the intermediate iterate $\boldsymbol{\theta}^{(k+1/2)}:=\pazocal{M}(\boldsymbol{\theta}^{(k)})$, which is required to reduce the loss, i.e., it has to satisfy the following condition: